OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2) A. 30 B. –15 C. no solution D. all real numbers is it c ?

OpenStudy (anonymous):

What's the question.

OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2)

OpenStudy (anonymous):

you're right.

OpenStudy (anonymous):

thank you !

Mehek (mehek14):

-15 + 45n = 45n - 30 15 + 45n = 45n no solution

OpenStudy (anonymous):

Determine which equation below has a solution of x = 1/2 A. 5x + 8 – 12x = –6x + 2x – 3 B. 3x – 1 – 4x= –2x + 7x – 4 C. –4x – 6 + x = –2x + x – 10 D. –3x – 8 + x = –4x + x – 10 this one ?

Mehek (mehek14):

solve for x for each one

OpenStudy (anonymous):

Can you re-post this question?

OpenStudy (anonymous):

okay which one ?

OpenStudy (anonymous):

hmm its not A

OpenStudy (anonymous):

not b

OpenStudy (anonymous):

d ?

Mehek (mehek14):

solve each one A 5x + 8 – 12x = –6x + 2x – 3 -7x + 8 = -4x - 3 -7x + 11 = -4x 11 = 3x x = 3.6 not A B 3x – 1 – 4x= –2x + 7x – 4 -x - 1 = 5x - 4 -x + 3 = 5x 3 = 6x x = 2 not B C –4x – 6 + x = –2x + x – 10 -3x - 6 = -x - 10 -2x = -4 x = 2 not C so ur left with D

OpenStudy (anonymous):

okay great ! thank you !!

Mehek (mehek14):

yw \(\ddot\smile\)